### v1

# 22-Split of Form Factors via BCFW Recursion Relation

Abstract
Recently, Cao:2025hio demonstrated the $2$-split for form factor under specific kinematic constraints. This factorization is analogous to that observed in scattering amplitudes. A key consequence of this structure is the presence of hidden zeros, where the form factors vanish on specific kinematic loci. We first establish these zeros and a new zero for the form factors of the composite operators ${\cal O}=\frac{1}{2}\mathrm{Tr}((\partial\phi)^{2})+\mathrm{Tr}(\phi^{3})$ and ${\cal O}=\mathrm{Tr}(F^{2})$, and then employ an inductive proof based on the BCFW recursion relation to prove the $2$-split factorization for any number of external particles.
1 Introduction
The study of scattering amplitudes has undergone a paradigm shift over the past decades. Recently, a remarkable new structure known as hidden zeros Arkani-Hamed:2023swr —the vanishing of various tree-level amplitudes at specific loci in kinematic space—has been discovered. This phenomenon has generated considerable interest and inspired extensive research from multiple perspectives Rodina:2024yfc ; Li:2024qfp ; Zhang:2024iun ; Zhou:2024ddy ; Zhang:2024efe ; Bartsch:2024amu ; Li:2025suo ; Huang:2025blb ; Backus:2025hpn ; De:2025bmf ; Jones:2025rbv . Closely related to this is a novel factorization property, the $2$-split Cao:2024gln ; Cao:2024qpp , which has likewise attracted substantial attention Arkani-Hamed:2024fyd ; Zhang:2024iun ; Zhou:2024ddy ; Zhang:2024efe ; Feng:2025ofq ; Feng:2025dci . The $2$-split arises when a particular constraint is imposed, under which the amplitude factorizes into two amputated currents.
An important question is whether these remarkable properties extend beyond scattering amplitudes to other fundamental observables. Form factors, which describe the coupling of an off-shell local operator to on-shell asymptotic states, provide a crucial bridge between on-shell and off-shell physics. They are defined as
$F_{n}^{{\cal O}}(1,2,\ldots,n;q)\equiv\int d^{D}x,e^{iq\cdot x}\braket{0|{\cal O}(x)|1,2,\ldots,n}\,$, (1)
where $q=\sum_{i=1}^{n}p_{i}$. Compared with general correlation functions, form factors are structurally simpler yet retain rich information about the operator. Understanding their analytic properties is therefore essential both for theoretical consistency and for phenomenological applications.
Indeed, the extension of the $2$-split to form factors has been demonstrated in Cao:2025hio , where it was shown that for the operators ${\cal O}=\tfrac{1}{2}\mathrm{Tr}[(\partial\phi)^{2}]+\mathrm{Tr}(\phi^{3})$ and ${\cal O}=\mathrm{Tr}(F^{2})$, string-theoretic integral representations exhibit the same factorization structure. Specifically, in certain kinematic regions, these form factors decompose into the product of a lower-point off-shell amplitude and a lower-point form factor.
The on-shell program has achieved substantial progress in the study of form factors Brandhuber:2010ad ; Bork:2010wf ; Brandhuber:2011tv ; Brandhuber:2012vm ; Brandhuber:2014ica ; Boels:2012ew ; Bork:2014eqa ; Frassek:2015rka ; Bork:2016hst ; Bork:2016xfn ; Yang:2016ear ; Bork:2017qyh ; Bianchi:2018peu ; Nandan:2018hqz ; Yang:2019vag ; Lin:2022jrp ; Lin:2023rwe . A cornerstone of this framework is the Britto–Cachazo–Feng–Witten (BCFW) recursion relation Britto:2004ap ; Britto:2005fq , which enables tree-level amplitudes to be constructed recursively from simpler, lower-point inputs. In this work, we extend the use of BCFW recursion relation to establish the $2$-split of tree-level form factors for specific composite operators in scalar and gauge theories. The proof closely parallels the amplitude case studied in Feng:2025ofq , but with a crucial difference: while the zeros of amplitudes naturally induce zeros for form factors, an additional hidden zero unique to form factors is required in order to establish the $2$-split.
The structure of this note is as follows. In Section 2, we identify the hidden zeros and a new zero relevant to the form factors of ${\cal O}=\tfrac{1}{2}\mathrm{Tr}[(\partial\phi)^{2}]+\mathrm{Tr}(\phi^{3})$ and ${\cal O}=\mathrm{Tr}(F^{2})$. In Section 3, we employ these results within an inductive proof based on BCFW recursion relation, thereby establishing the $2$-split factorization for arbitrary multiplicity. We conclude in Section 4 with a brief summary and outlook.
2 Hidden zeros
Let us consider the operator ${\cal O}={\cal L}_{\phi}\equiv\frac{1}{2}\rm Tr\left((\partial\phi)^{2}\right)+\rm Tr(\phi^{3})$ with external particles taken to be scalars in the $\mathrm{Tr}(\phi^{3})$ theory. We select three distinguished particles $i,j,k$ and partition the remaining external legs into two disjoint sets $A$ and $B$, such that $A\cup B=\{1,\ldots,n\}\setminus\{i,j,k\}$. We then impose the constraint
$s_{a,b}=s_{a,q}=0\,,\quad\forall a\in A,b\in B\,$, (2)
where $s_{a,b}=2p_{a}\cdot p_{b}$ and $s_{a,q}=2p_{a}\cdot q$. Under these kinematic conditions, the form factor $F^{{\cal L}_{\phi}}_{n}(1,\ldots,n;q)$ exhibits a $2$-split factorization:
$F^{{\cal L}_{\phi}}_{n}(i,A,j,B(k);q)\to{\cal A}^{\phi^{3}}_{|A|+3}(i,A,j;\kappa)F^{{\cal L}_{\phi}}_{|B|+3}(j,B,i;\kappa^{\prime},q)\,$, (3)
where $B(k)\equiv B\cup\{k\}$ preserves the ordering [^1] For instance, $B=\{1,2,4\},k=3$, the $B(k)=\{1,2,3,4\}$.. We use the semicolon to single out the off-shell leg, which is not written according to the ordering. $\kappa,\kappa^{\prime}$ take the position of that of the leg kk [^2] For example, $k=n$, then $(i,A,j;\kappa)=(i,A,j,\kappa)$ and $(j,B,j;\kappa^{\prime})=(j,\ldots,n-1,\kappa^{\prime},1,\ldots,i)$..
The resulting amplitude is known to exhibit hidden zeros Arkani-Hamed:2023swr , namely
${\cal A}^{\phi^{3}}_{|A|+3}(i,A,j;\kappa)\xrightarrow{s_{a,\kappa}=0}0\,$, (4)
where momentum conservation implies $p_{\kappa}=p_{B}+p_{k}-q$, so that $s_{a,\kappa}=s_{a,k}$. Consequently, the form factor also vanishes under the same conditions:
[…]
In addition, the form factor possesses other hidden zeros not directly inherited from the amplitude. Specifically, we claim that
$F^{{\cal L}_{\phi}}_{n}(i,\alpha,j;q)\xrightarrow{s_{a,q}=0,\forall a\in\alpha}0\,.$ (6)
This property can be proven inductively. The base case is the three-point form factor,
$F^{{\cal L}_{\phi}}_{3}(1,2,3;q)=\frac{(s_{1,2}+s_{2,3})(s_{1,2}+s_{1,3})(s_{1,3}+s_{2,3})}{(s_{1,2}s_{1,3}s_{2,3})}\,.$ (7)
This expression clearly vanishes when the condition $s_{2,q}=s_{1,2}+s_{2,3}=0$ is applied.
Assuming (6) holds for form factors with fewer than $n$ points, we can establish it for $F^{{\cal L}_{\phi}}_{n}$ using the BCFW recursion relation. First, we must confirm that the form factor has no boundary contribution under a BCFW shift:
$p_{i}\to\hat{p}_{i}\to p_{i}+zr,~~~~~~p_{j}\to\hat{p}_{j}\to p_{j}+zr$ (8)
with $r^{2}=r\cdot p_{i}=r\cdot p_{j}=0$. The primary difference between this form factor and a pure amplitude arises from the kinematic term $\frac{1}{2}\rm Tr\left((\partial\phi)^{2}\right)$ in the operator ${\cal L}_{\phi}$. This introduces a non-trivial numerator $p\cdot q$, but it also provides an additional propagator. The net scaling at large $z$ is therefore favorable, allowing the BCFW recursion relation to be applied in the form factor Brandhuber:2010ad :
$F^{{\cal L}_{\phi}}_{n}(i,\alpha,j;q)=\sum_{I}F^{{\cal L}_{\phi}}_{L}(\hat{i},\alpha_{L},\hat{I};q)\frac{1}{P_{I}^{2}}{\cal A}^{\phi^{3}}_{R}(-\hat{I},\alpha_{R},\hat{j})+{\cal A}^{\phi^{3}}_{L}(\hat{i},\alpha_{L},\hat{I})\frac{1}{P_{I}^{2}}F^{{\cal L}_{\phi}}_{R}(-\hat{I},\alpha_{R},\hat{j};q)\,.$ (9)
In the kinematic limit where $s_{a,q}=0$ for all $a\in\alpha$, every term in the BCFW sum involves a lower-point form factor $F^{{\cal L}_{\phi}}_{L}$ or $F^{{\cal L}_{\phi}}_{R}$ evaluated on a subset of these legs. By the induction hypothesis, these lower-point form factors vanish. Consequently, the entire right-hand side is zero.
This completes the proof, confirming that (6) is valid for all $n\geqslant 3$.
Similarly, the Yang-Mills (YM) form factor for the operator ${\cal O}=\rm Tr(F^{2})$ possesses hidden zeros. The YM form factor also exhibits a $2$-split factorization on a different kinematic locus:
$F^{\rm Tr(F^{2})}_{n}(i,A,j,B(k);q)\,\xrightarrow[]{}\,{\cal J}^{\rm mixed}_{|A|+3}(i^{\phi},A,j^{\phi};\kappa^{\phi})\,F^{{\rm Tr}(F^{2})}_{|B|+3}(j,B,i;\kappa^{\prime},q)_{\mu}\epsilon_{k}^{\mu}\,$, (10)
under the kinematic constraint of (2) and the polarization constraint:
$\epsilon_{a}\cdot\epsilon_{b^{\prime}}=\epsilon_{a}\cdot p_{b}=\epsilon_{b^{\prime}}\cdot p_{a}=\epsilon_{a}\cdot q=0\,,\quad\forall a\in A,b\in B,b^{\prime}\in B\cup\{i,j,k\}\,.$ (11)
Here, ${\cal J}^{\rm mixed}_{|A|+3}$ is a current with $|A|$ gluons and 33 $\phi$’s. The resulting current is known to have hidden zeros Arkani-Hamed:2023swr , vanishing when $s_{a,\kappa}=\epsilon_{a}\cdot p_{\kappa}=0$. Given the momentum definitions, this condition becomes $s_{a,k}=\epsilon_{a}\cdot p_{k}=0$. This implies that the entire form factor vanishes on this locus:
${\cal J}^{\rm mixed}_{|A|+3}(i^{\phi},A,j^{\phi};\kappa^{\phi})\xrightarrow{s_{a,\kappa}=\epsilon_{a}\cdot p_{\kappa}=0}0\,$, (12)
Thus, the YM form factor vanishes with the same constraint:
[…]
We claim
$F^{{\cal L}_{\phi}}_{n}(i,\alpha,j;q)\xrightarrow{s_{a,q}=\epsilon_{a}\cdot q=0,\forall a\in\alpha}0\,.$ (14)
We can prove it by induction. The simplest case is $F^{\mathrm{Tr}(F^{2})}_{3}(1,2,3;q)$, given in Dong:2022bta :
$F^{\mathrm{Tr}(F^{2})}_{3}(1,2,3;q)=\frac{1}{s_{1,2}}\left({\rm tr}(2,3)p_{2}\cdot\epsilon_{1}-{\rm tr}(1,3)p_{1}\cdot\epsilon_{2}+2{\rm tr}(1,2,3)\right)+{\rm cylic}(1,2,3)\,$, (15)
where ${\rm tr}(\alpha_{1},\ldots,\alpha_{m})\equiv(f_{\alpha_{1}})_{\mu_{1}}^{\mu_{2}}\cdots(f_{\alpha_{m}})_{\mu_{m}}^{\mu_{1}}$ and $f_{i}^{\mu\nu}=p_{i}^{\mu}\epsilon_{i}^{\nu}-p_{i}^{\mu}\epsilon_{i}^{\nu}$. This expression can be verified to vanish under the conditions $s_{2,q}=s_{1,2}+s_{2,3}=0$ and $\epsilon_{2}\cdot q=\epsilon_{2}\cdot p_{1}+\epsilon_{2}\cdot p_{3}=0$.
The BCFW recursion relation for the YM form factor is Brandhuber:2010ad
[…]
$+{\cal A}^{\phi^{3}}_{L}(\hat{i},\alpha_{L},\hat{I})\frac{1}{P_{I}^{2}}F^{\mathrm{Tr}(F^{2})}_{R}(-\hat{I},\alpha_{R},j;q)\Big{)}\,$, (16)
where the $\sum_{h}$ is sum over the helicities of the internal gluon $I$. As before, by the induction hypothesis, the lower-point form factors on the right-hand side, $F^{\mathrm{Tr}(F^{2})}_{L}$ and $F^{\mathrm{Tr}(F^{2})}_{R}$, vanish in the specified kinematic limit. Thus, (14) holds for all $n\geqslant 3$.
3 $2$-split
3.1 ${\cal O}=\frac{1}{2}\rm Tr\left((\partial\phi)^{2}\right)+\rm Tr(\phi^{3})$
We will now prove the 2-split behavior of the form factor, (3), by induction on the number of particles. The base case, where the set $A$ is empty ($|A|=0$), is trivial, as the three-point amplitude ${\cal A}^{\phi^{3}}_{3}(i,j;\kappa)=1$. For the inductive hypothesis, we assume that the $2$-split property, and therefore the associated hidden zero under the constraint of (2), holds for all form factors $F^{{\cal L}_{\phi}}_{m}$ with fewer than $n$ particles ($m<n$). We prove the property for $F^{{\cal L}_{\phi}}_{n}$ using the BCFW recursion relation. Under the $2$-split kinematic constraints, the inductive hypothesis implies that most BCFW diagrams vanish. Only the four diagram types shown in Fig. 1 yield non-zero contributions. Each diagram consists of two terms: one where the form factor operator is on the left sub-diagram and the amplitude is on the right, and another with the roles reversed.
Figure 1: The four non-vanishing BCFW diagrams that contribute to the form factors.
The first diagram: In this diagram, the BCFW channel splits the set of particles $A$ into $A_{1}$ and $\bar{A}_{1}$.The contribution is:
$\sum_{I_{1}}F^{\cal{L}_{\phi}}(\hat{i},A_{1},\hat{I}_{1};q)\frac{1}{P_{I_{1}}^{2}}{\cal A}^{\phi^{3}}(-\hat{I}_{1},\bar{A_{1}},\hat{j},B;k)+{\cal A}^{\phi^{3}}(\hat{i},A_{1},\hat{I}_{1})\frac{1}{P_{I_{1}}^{2}}F^{\cal{L}_{\phi}}(-\hat{I}_{1},\bar{A_{1}},\hat{j},B;k,q)\,$, (17)
where first term vanishes because the lower-point form factor $F^{\cal{L}_{\phi}}(\hat{i},A_{1},\hat{I}_{1};q)$ is subject to a hidden zero condition from (14). The second term can be analyzed by applying the $2$-split property to the lower-point form factor inside the sum:
$\sum_{I_{1}}{\cal A}^{\phi^{3}}(\hat{i},A_{1},\hat{I}_{1})\frac{1}{P_{I_{1}}^{2}}\Big{(}{\cal A}^{\phi^{3}}(-\hat{I}_{1},\bar{A_{1}},\hat{j};\kappa_{I_{1}})F_{|B|+3}^{\cal{L}_{\phi}}(\hat{j},B,-\hat{I}_{1};\kappa_{I_{1}}^{\prime},q)\Big{)}$
$=$ $\left(\sum_{I_{1}}{\cal A}^{\phi^{3}}(\hat{i},A_{1},\hat{I}_{1})\frac{1}{P_{I_{1}}^{2}}{\cal A}^{\phi^{3}}(-\hat{I}_{1},\bar{A_{1}},\hat{j};\kappa)\right)F_{|B|+3}^{{\cal L}_{\phi}}(\hat{j},B,\hat{i};\kappa^{\prime},q)\,.$ (18)
In the first line, the momentum conservation determines
$p_{\kappa_{I_{1}}^{\prime}}=p_{k}+p_{\bar{A_{1}}},~~~~p_{\kappa_{I_{1}}}=p_{k}+p_{B}-q=p_{\kappa}\,.$ (19)
In the second line, we have used a key relation:
$F_{|B|+3}^{{\cal L}_{\phi}}(\hat{j},B,-\hat{I}_{1};\kappa_{I_{1}}^{\prime},q)=F_{|B|+3}^{{\cal L}_{\phi}}(\hat{j},B,\hat{i};\kappa^{\prime},q)\,.$ (20)
This equality holds because the propagator along the line connecting $\hat{I}_{1},\hat{j}$ on the left-hand side, dot products with the shifted momentum $\hat{P}_{I_{1}}$ reduce to dot products with $\hat{p}_{i}$:
$(-\hat{P}_{I_{1}}+p_{B^{\prime}})^{2}$ $=$ $p_{B^{\prime}}^{2}-2\hat{P}_{I_{1}}\cdot p_{B^{\prime}}=p_{B^{\prime}}^{2}+2(\hat{p}_{i}+p_{A_{1}})\cdot p_{B^{\prime}}$ (21)
$=$ $p_{B^{\prime}}^{2}+2\hat{p}_{i}\cdot p_{B^{\prime}}=(\hat{p}_{i}+p_{B^{\prime}})^{2}\,.$
Here, we have used the condition (2) and $\hat{p}_{i}^{2}=\hat{P}_{I_{1}}^{2}=0$. For the block containing the leg $\kappa_{I_{1}}^{\prime}$ on the left-hand side, the inverse of massive propagators is given by
$(p_{\kappa_{I}^{\prime}}+p_{B_{i}})^{2}-(p_{\kappa_{I}^{\prime}})^{2}$ $=$ $p_{B_{i}}^{2}+2p_{B_{i}}\cdot p_{\kappa_{I}^{\prime}}=p_{B_{i}}^{2}+2p_{B_{i}}\cdot(p_{k}+p_{A_{1R}})$ (22)
$=$ $p_{B_{i}}^{2}+2p_{B_{i}}\cdot(p_{k}+p_{A})=(p_{\kappa^{\prime}}+p_{B_{i}})^{2}-(p_{\kappa^{\prime}})^{2}$
Similarly, the numerator factor from the operator $\frac{1}{2}\rm Tr\left((\partial\phi)^{2}\right)$ is same on the both-hand sides of (20):
$q\cdot p_{\kappa_{I_{1}}^{\prime}}=q\cdot(p_{k}+p_{\bar{A_{1}}})=q\cdot p_{k}\,,\quad-q\cdot\hat{p}_{I_{1}}=q\cdot(\hat{p}_{i}+p_{A_{1}})=q\cdot\hat{p}_{i}\,.$ (23)
A similar analysis applies to the other three diagrams.
  • The second diagram:
[…]
    $=$ $\left(\sum_{I_{2}}{\cal A}^{\phi^{3}}(\hat{i},A_{2},\hat{I}_{2};\kappa)\frac{1}{P_{I_{2}}^{2}}{\cal A}^{\phi^{3}}(-\hat{I}_{2},\bar{A_{2}},\hat{j})\right)F_{|B|+3}^{{\cal L}_{\phi}}(\hat{j},B(\kappa^{\prime}),\hat{i};q)\,.$ (24)
  • The third diagram:
[…]
    $=$ ${\cal A}^{\phi^{3}}(\hat{i},A,\hat{j},B;\kappa)\Big{(}\sum_{I_{3}}F^{{\cal L}_{\phi}}(B_{3},\hat{i},\hat{I}_{3};q)\frac{1}{P_{I_{3}}^{2}}{\cal A}^{\phi^{3}}(-\hat{I}_{3},\hat{j},\bar{B}_{3};\kappa^{\prime})$
    $~~~~~~~+{\cal A}^{\phi^{3}}(B_{3},\hat{i},\hat{I}_{3})\frac{1}{P_{I_{3}}^{2}}F^{{\cal L}_{\phi}}(-\hat{I}_{3},\hat{j},\bar{B}_{3};\kappa^{\prime},q)\Big{)}\,.$ (25)
  • The forth diagram:
[…]
    $={\cal A}^{\phi^{3}}(\hat{i},A,\hat{j},B;\kappa)\Big{(}\sum_{I_{4}}F^{{\cal L}_{\phi}}(B_{4},\hat{i},\hat{I}_{4};\kappa^{\prime},q)\frac{1}{P_{I_{4}}^{2}}{\cal A}^{\phi^{3}}(-\hat{I}_{4},A,\hat{j},\bar{B}_{4})$
    $~~~~~~~+{\cal A}^{\phi^{3}}(B_{4},\hat{i},\hat{I}_{4};\kappa^{\prime})\frac{1}{P_{I_{4}}^{2}}F^{{\cal L}_{\phi}}(-\hat{I}_{4},\hat{j},\bar{B}_{4};q)\Big{)}\,.$ (26)
Summing the contributions from these four non-vanishing diagrams in (17), (3.1), (3.1), and (3.1) precisely reconstructs the BCFW expansion for the right-hand side of the $2$-split formula, (3). The first 2 diagrams build the right-hand current, while the last 2 diagrams build the right-hand form factor.
This completes the inductive proof for the $2$-split of the form factor $F^{{\cal L}_{\phi}}_{n}$ (3).
3.2 ${\cal O}=\rm Tr(F^{2})$
The method to prove the $2$-split for YM form factor is analogous to the scalar case, relying on induction and the BCFW recursion relation. For the set $A$ is empty ($|A|=0$), the case is trivial, as the three-point current ${\cal J}^{\rm mixed}_{3}(i^{\phi},j^{\phi};\kappa^{\phi})=1$. For the inductive hypothesis, we assume that the $2$-split, and therefore the associated hidden zeros (13), holds for all form factors with fewer than $n$ external particles. We prove the property for $F^{\mathrm{Tr}(F^{2})}_{n}$ using the BCFW recursion relation. We apply a BCFW shift to legs $i$ and $j$, choosing the shifted polarizations $\hat{\epsilon}_{i}$ and $\hat{\epsilon}_{j}$ to satisfy the constraints of (11). Due to the inductive hypothesis, only the four BCFW diagram types previously shown in Fig. 1 give non-zero contributions.
Consider the first diagram, the contribution is:
$\sum_{I_{1}}\sum_{h}{\cal A}^{\rm YM}(\hat{i},A_{1},\hat{I}_{1}){F^{\mathrm{Tr}(F^{2})}}(-\hat{I}_{1},\bar{A}_{1},\hat{j},B(k);q)\,.$ (27)
Here, the term where the form factor is on the left sub-diagram, $F^{\mathrm{Tr}(F^{2})}(\hat{i},A_{1},\hat{I}_{1})$, vanishes due to the polarization constraints in (11), so it has been omitted.
A key step is to resolve the polarization sum. The polarization constraints (11) require that the shifted polarization $\hat{\epsilon}_{i}$ can only contract with the polarization of the internal gluon, $\epsilon_{I}^{h}$. otherwise the amplitude ${\cal A}^{\rm YM}(\hat{i},A_{1},\hat{I}_{1})$ is annihilated due to the condition (11). Using $\sum_{h}\,\big{(}\epsilon_{I_{1}}^{h}\,\epsilon_{-I_{1}}^{-h}\big{)}^{\mu\nu}\,\sim\,g^{\mu\nu}$ we can effectively absorb $\hat{\epsilon}_{i}$ into the form factor by replacing $\epsilon_{-I_{1}}^{-h}$ with $\hat{\epsilon}_{i}$. This eliminates the helicity sum and converts the YM amplitude into a YM-Scalar (YMS) amplitude, where legs $\hat{i}$ and $\hat{I}$ are treated as scalars.
Now, applying the inductive hypothesis to the lower-point form factor, we get:
$F^{\mathrm{Tr}(F^{2})}(-\hat{I}^{\hat{\epsilon}_{i}},\bar{A}_{1},\hat{j},B(k);q)\xrightarrow[]{}\,{\cal J}^{\rm mixed}(-\hat{I}^{\phi},\bar{A}_{1},\hat{j}^{\phi};\kappa_{I_{1}}^{\phi})F^{\mathrm{Tr}(F^{2})}(\hat{j},B,\hat{I}^{\hat{\epsilon}_{i}};\kappa^{\prime}_{I_{1}},q)_{\mu}\epsilon_{k}^{\mu}\,$, (28)
where the superscript $\hat{\epsilon}_{i}$ is to emphasize that the polarization of the particle $\hat{I}$ is $\hat{\epsilon}_{i}$. Thus we arrive at
$\sum_{I_{1}}\sum_{h}\,{\cal A}^{\rm YM}(\hat{i},A_{1},\hat{I}_{1})\,{1\over P_{I_{1}}^{2}}\,F^{\mathrm{Tr}(F^{2})}(-I_{1},\bar{A}_{1},\hat{j},B(k);q)$
$=$ $\sum_{I_{1}}{\cal A}^{\rm YMS}(\hat{i}^{\phi},A_{1},\hat{I}_{1}^{\phi})\,{1\over P_{I_{1}}^{2}}\,{\cal J}^{\rm mixed}(-\hat{I}_{1}^{\phi},\bar{A}_{1},\hat{j}^{\phi};\kappa_{I_{1}}^{\phi})F^{\mathrm{Tr}(F^{2})}(\hat{j},B,\hat{I}_{1}^{\hat{\epsilon}_{i}};\kappa^{\prime}_{I_{1}},q)_{\mu}\epsilon_{k}^{\mu}$
$=$ $\sum_{I_{1}}{\cal A}^{\rm YMS}(\hat{i}^{\phi},A_{1},\hat{I}_{1}^{\phi})\,{1\over P_{I_{1}}^{2}}\,{\cal J}^{\rm mixed}(-\hat{I}_{1}^{\phi},\bar{A}_{1},\hat{j}^{\phi};\kappa^{\phi})F^{\mathrm{Tr}(F^{2})}(\hat{j},B,\hat{i};\kappa^{\prime},q)_{\mu}\epsilon_{k}^{\mu}\,$, (29)
where $p_{\kappa_{I_{1}}}=p_{B}+p_{k}-q=p_{\kappa}$, $p_{\kappa^{\prime}_{I_{1}}}=p_{k}+p_{\bar{A_{1}}}$. To get the third equation, we need to prove
$F^{\mathrm{Tr}(F^{2})}(\hat{j},B,\hat{I}_{1}^{\hat{\epsilon}_{i}};\kappa^{\prime}_{I_{1}},q)_{\mu}\epsilon_{k}^{\mu}\overset{?}{=}F^{\mathrm{Tr}(F^{2})}(\hat{j},B,\hat{i};\kappa^{\prime},q)_{\mu}\epsilon_{k}^{\mu}$ (30)
Their denominators are identical, as argued in the scalar case (20). The numerator is a function of Lorentz contraction. The non-kinematic part from the $\mathrm{Tr}(F^{2})$ operator is the same, as argued in Feng:2025ofq . The vertex inserted by the kinematic term is $V^{\mu\nu}=p\cdot qg^{\mu\nu}-p^{\mu}q^{\nu}$. Since the polarizations are same in both sides of (30) and $q\cdot\{\hat{p}_{I_{1}},p_{\kappa^{\prime}_{I_{1}}}\}=q\cdot\{\hat{p}_{i},p_{\kappa^{\prime}}\}$. Thus, (30) is correct. Analogously, the second diagram is
[…]
$=$ $\sum_{I_{2}}{\cal J}^{\rm mixed}(\hat{i}^{\phi},A_{2},\hat{I}_{2}^{\phi};\kappa^{\phi})\,{1\over P_{I_{2}}^{2}}\,{\cal A}^{\rm YM}(-I_{2}^{\phi},\bar{A}_{2},\hat{j}^{\phi})F^{\mathrm{Tr}(F^{2})}(\hat{j},B,\hat{i};\kappa^{\prime},q)_{\mu}\epsilon_{k}^{\mu}\,.$ (31)
Together, these two diagrams build the right-hand current in the 2-split formula.
For the remaining diagrams, a fully rigorous proof is complex. Instead, we demonstrate the factorization by considering a specific kinematic configuration that satisfies all constraints. We decompose the full kinematic space into two orthogonal subspaces, ${\cal S}={\cal S}_{1}\oplus{\cal S}_{2}$. Momenta and polarizations of legs in set $A$ lie in ${\cal S}_{1}$, while those for legs in set $B$, along with $q$ and polarizations $\epsilon_{i,j,k}$ lie in ${\cal S}_{2}$.
Consider the third diagram’s contribution:
$\sum_{I_{3}}\sum_{h}$ $\Big{(}F^{\mathrm{Tr}(F^{2})}(\hat{i},B_{3},\hat{I}_{3};q)\frac{1}{P_{I_{3}}^{2}}{\cal A}^{\rm YM}(-\hat{I}_{3},A,\hat{j},\bar{B}_{3}(k))$
$+{\cal A}^{\rm YM}(\hat{i},B_{3},\hat{I}_{3})\frac{1}{P_{I_{3}}^{2}}F^{\mathrm{Tr}(F^{2})}(-\hat{I}_{3},A,\hat{j},\bar{B}_{3}(k);q)\Big{)}\,.$ (32)
For the amplitude ${\cal A}^{\rm YM}(\hat{i},B_{3},\hat{I}_{3})$ to be non-zero, the effective polarization $\epsilon_{I_{3}}^{H}$ must lie in ${\cal S}_{1}$. This is because all other momenta and polarizations in this amplitude lie in ${\cal S}_{1}$, except $\hat{p}_{i}$, lie in ${\cal S}_{1}$. For $\hat{p}_{i}$, we can use the momentum conservation $\hat{p}_{i}=-\hat{P}_{I_{3}}-p_{B_{3}}$ and the on-shell condition $\epsilon_{I_{3}}^{H}\cdot\hat{P}_{I_{3}}=0$, to confirm $\epsilon^{H}_{I_{3}}\cdot\hat{p}_{i}=0$. Thus, any component of $\epsilon_{I_{3}}^{H}$ in ${\cal S}_{2}$ would lead to a vanishing contraction.
This constraint forces the factorization of the sub-diagrams according to the 2-split hypothesis. The amplitude and the form factor on the right-hand side both split, factoring out the same lower-point current:
$F^{\mathrm{Tr}(F^{2})}(-\hat{I}_{3},A,\hat{j},\bar{B}_{3}(k);q)\xrightarrow[]{}\,{\cal J}^{\rm mixed}(-\hat{I}_{3}^{\phi},A,\hat{j}^{\phi};\kappa^{\phi}_{I_{3}}){F}^{\mathrm{Tr}(F^{2})}(\hat{j},\bar{B}_{3},-\hat{I}_{3};\kappa^{\prime}_{I_{3}},q)_{\mu}\epsilon^{\mu}_{k}\,$,
${\cal A}^{\rm YM}(-\hat{I}_{3},A,\hat{j},\bar{B}_{3}(k))\xrightarrow[]{}\,{\cal J}^{\rm mixed}(-\hat{I}_{3}^{\phi},A,\hat{j}^{\phi};\kappa^{\phi}_{I_{3}}){\cal J}^{\rm YM}(\hat{j},\bar{B}_{3},-\hat{I}_{3};\kappa^{\prime}_{I_{3}})\,.$ (33)
Here, $\kappa^{\prime}_{I_{3}}=\kappa^{\prime}$ follow from momentum conservation. From the scalar case discussion and Feng:2025ofq , we know
${\cal J}^{\rm mixed}(-\hat{I}_{3}^{\phi},A,\hat{j}^{\phi};\kappa^{\phi}_{I_{3}})={\cal J}^{\rm mixed}(i^{\phi},A,\hat{j}^{\phi};\kappa^{\phi})$ (34)
Plugging (3.2), and (34) into (3.2), we can factor out this common current:
${\cal J}^{\rm mixed}(i^{\phi},A,\hat{j}^{\phi};\kappa^{\phi})\sum_{I_{3}}\sum_{h}$ $\Big{(}F^{\mathrm{Tr}(F^{2})}(\hat{i},B_{3},\hat{I}_{3};q)\frac{1}{P_{I_{3}}^{2}}{\cal J}^{\rm YM}(-\hat{I}_{3},\hat{j},\bar{B}_{3};\kappa^{\prime})$
$+{\cal A}^{\rm YM}(\hat{i},B_{3},\hat{I}_{3})\frac{1}{P_{I_{3}}^{2}}F^{\mathrm{Tr}(F^{2})}(-\hat{I}_{3},\hat{j},\bar{B}_{3};\kappa^{\prime},q)\Big{)}$ (35)
The same logic applies to the fourth diagram:
${\cal J}^{\rm mixed}(i^{\phi},A,\hat{j}^{\phi};\kappa^{\phi})\sum_{I_{3}}\sum_{h}$ $\Big{(}F^{\mathrm{Tr}(F^{2})}(\hat{i},B_{3},\hat{I}_{3};\kappa^{\prime},q)\frac{1}{P_{I_{3}}^{2}}{\cal J}^{\rm YM}(-\hat{I}_{3},\hat{j},\bar{B}_{3})$
$+{\cal A}^{\rm YM}(\hat{i},B_{3},\hat{I}_{3};\kappa^{\prime})\frac{1}{P_{I_{3}}^{2}}F^{\mathrm{Tr}(F^{2})}(-\hat{I}_{3},\hat{j},\bar{B}_{3};q)\Big{)}$ (36)
The sums in the brackets for the third and forth diagrams then combine to form the BCFW construction of the right-hand form factor.
Therefore, combining Eqs. (3.2), (3.2), (3.2), and (3.2) yields precisely the BCFW recursion relation for the right-hand side of (10).
4 Conclusion
In this work, our central result is the proof of the $2$-split factorization for the form factors of the operators ${\cal O}=\tfrac{1}{2}\mathrm{Tr}((\partial\phi)^{2})+\mathrm{Tr}(\phi^{3})$ and ${\cal O}=\mathrm{Tr}(F^{2})$. By imposing specific kinematic constraints that partition the external legs, Cao:2025hio showed that an $n$-point form factor factorizes into a product of a lower-point amplitude and a lower-point form factor. We prove this behavior by induction. The base cases were established directly, and the inductive step was carried out by applying the BCFW recursion relation. We showed that under the $2$-split kinematic conditions, only a select few BCFW diagrams contribute, and their sum precisely reconstructs the factorized form. A crucial element of our analysis was the identification of hidden zeros, which are necessary for the consistency of the inductive proof. Unlike the proof for amplitudes in Feng:2025ofq , these hidden zeros are insufficient to establish the $2$-split property for form factors. An additional zero is required, ensuring the vanishing of $F(\hat{i},A,\hat{I};q)$ (and $F(\hat{I},A,\hat{j};q)$).
The techniques and results presented here further bridge the gap between the study of scattering amplitudes and form factors, highlighting the universal applicability of on-shell methods. This opens several avenues for future research. An obvious next step would be to investigate whether similar factorization properties hold for form factors of operators in effective field theories or theories with massive particles would also be a valuable direction. Ultimately, it remains an open question whether the form factor can be bootstrapped from its zeros, as in Rodina:2024yfc .
Acknowledgments
We thank Qu Cao for valuable discussions, and Bo Feng and Kang Zhou for helpful comments on the manuscript. This work was supported by the CSRC Postdoctoral Budget.